OpenStudy (anonymous):

The vertex of this parabola is at (-2, 5). Which of the following could be its equation?

rebeccaxhawaii (rebeccaxhawaii):

choices?

OpenStudy (anonymous):

c

rebeccaxhawaii (rebeccaxhawaii):

please do not give direct answers

OpenStudy (anonymous):

can u teach me how to do it

rebeccaxhawaii (rebeccaxhawaii):

i would suggest graphing it https://desmos.com

OpenStudy (anonymous):

the site is blocked

rebeccaxhawaii (rebeccaxhawaii):

https://www.desmos.com/calculator

OpenStudy (anonymous):

plug in x, if the value of y matches that is your answer

OpenStudy (anonymous):

how do i plug in x?

rebeccaxhawaii (rebeccaxhawaii):



OpenStudy (anonymous):

y = c(x-x1) + y1 where (x1, y1) is vertex that is (-2,5) in this case... c can be any number which is 3 in this case

OpenStudy (mathmale):

"The vertex of this parabola is at (-2, 5). Which of the following could be its equation? " Think: the "vertex form of the equation of a parabola" is y=a(x-h)^2+k, where (h,k) represents the vertex. Take the given vertex and subst. its coordinates into this "vertex form." Try it. Your result should eliminate all but 1 or 2 of the possible answer choices.

OpenStudy (anonymous):

where did the 3 come from?

OpenStudy (anonymous):

it doest have to be 3 , it can be any number

OpenStudy (anonymous):

so that number is random?

OpenStudy (mathmale):

You don't have enuf info to determine that. You're not asked to find the y-intercept, but rather the answer form closest enough to the vertex form of the eq'n of a parabola.

OpenStudy (anonymous):

depends how wide or thinner you want your parabola to be...in the question above, you dont need to care about such thing but only vertex so, it can be anything

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

so the second part would be (x-2)²+5

OpenStudy (mathmale):

Check your (x-2). Not correct.

OpenStudy (anonymous):

why would it not be x-2?

OpenStudy (mathmale):

Vertex is at (-2,5), mpt at (2,5).

OpenStudy (anonymous):

what is mpt

OpenStudy (mathmale):

Review the eqution I gave you: y=a(x-h)^2+k, and the given vertex: (-2,5)

OpenStudy (mathmale):

mpt=not. Sorry.

OpenStudy (anonymous):

why would it be x+2

OpenStudy (mathmale):

because you are forming (x-h) from h=-2. Try it.

OpenStudy (mathmale):

(x-2): incorrect (x-[-2]) = (x+2): correct

OpenStudy (anonymous):

oh -+- = +

OpenStudy (mathmale):

Yes.

OpenStudy (anonymous):

AnswerC